# SteerDiff: Steering towards Safe Text-to-Image Diffusion Models

### v1

@@
Abstract
Text-to-image (T2I) diffusion models have drawn attention for their ability to generate high-quality images with precise text alignment. However, these models can also be misused to produce inappropriate content. Existing safety measures, which typically rely on text classifiers or ControlNet-like approaches, are often insufficient. Traditional text classifiers rely on large-scale labeled datasets and can be easily bypassed by rephrasing. As diffusion models continue to scale, fine-tuning these safeguards becomes increasingly challenging and lacks flexibility. Recent red-teaming attack researches further underscore the need for a new paradigm to prevent the generation of inappropriate content. In this paper, we introduce SteerDiff, a lightweight adaptor module designed to act as an intermediary between user input and the diffusion model, ensuring that generated images adhere to ethical and safety standards with little to no impact on usability. SteerDiff identifies and manipulates inappropriate concepts within the text embedding space to guide the model away from harmful outputs. We conduct extensive experiments across various concept unlearning tasks to evaluate the effectiveness of our approach. Furthermore, we benchmark SteerDiff against multiple red-teaming strategies to assess its robustness. Finally, we explore the potential of SteerDiff for concept forgetting tasks, demonstrating its versatility in text-conditioned image generation.
@@ Abstract
1 Introduction
@@
Text-to-image (T2I) diffusion models have attracted attention for their out-of-the-box functionality and the superior quality of their generated images. Using models like Stable Diffusion (Rombach et al., 2022; 2021) or DALL-E (Ramesh et al., 2022), users can use simple natural language descriptions as input to generate high-quality images with precise text alignment. This capability is largely contributed by pre-trained language models (Hammoud et al., 2024; Jin et al., 2020; Dosovitskiy, 2020) that learn and reflect the underlying syntax and semantics, as well as by extensive multimodal training datasets that encompass a wide range of text-to-image aligned content. However, these training methods also introduce risks of generating inappropriate images to the models. Thus, preventing the generation of inappropriate images is both critical and urgent.
To this end, existing T2I diffusion models have integrated several safety strategies to prevent inappropriate content generation. Stable diffusion incorporates a Not Safe For Work (NSFW) post-generation filter and ad-hoc filter during training to avoid generating unsafe images (Schuhmann et al., 2022; Schramowski et al., 2023; Rando et al., 2022). However, existing text-based attacks (Li et al., 2018; Jin et al., 2020; Garg & Ramakrishnan, 2020; Maus et al., 2023) can mislead the classification mechanisms in the filter by rephrasing “A photo of a billboard showing a naked man.” into “A photo of a billboard showing an LGBT man in an explicit position”. Additionally, users report that removing explicit images and other subjects from training data may have had a negative impact on the output quality, thus harming the utility of the models (Rombach et al., 2022).
The prevention of inappropriate content generation in AI models has primarily been approached through concept removal, which aims to control the generation process itself. Two primary strategies have been explored: (1) using ControlNet-like (Zhang et al., 2023a) structures to guide the diffusion process (Schramowski et al., 2023; Rando et al., 2022), and (2) identifying and pruning weights within the diffusion model (Gandikota et al., 2023). However, these methods face significant challenges as generative models grow in scale and complexity. The increasing size of models makes these approaches computationally expensive and impractical (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022). Additionally, simply guiding the generative process introduces vulnerabilities, leaving models open to jailbreaking attempts (Zhang et al., 2023b; Qu et al., 2023). These limitations highlight the need for more efficient and robust solutions.
@@
Although the aforementioned safety mechanisms have shown effectiveness according to their respective evaluation schemes, recent red-teaming studies demonstrate their potential flaws (Zhang et al., 2023b; Qu et al., 2023). Chin et al. (2023) shows that approximately half of the prompts, which were originally mitigated by existing safety mechanisms, can be manipulated by their Prompting4Debugging (P4D) to become problematic. Similarly, Unsafe Diffusion (Qu et al., 2023) found that 14.56% of generated images across four state-of-the-art T2I models and their four prompt datasets were unsafe, underscoring the vulnerability of these models to generating harmful content. Moreover, black-box jail-breaking approaches Jailbreak Prompt Attack (JPA) and SneakyPrompt (Ma et al., 2024; Yang et al., 2024) successfully attack both online services and offline T2I models with the current safety mechanisms.
@@
In this work, we propose SteerDiff, a two-stage lightweight adaptor model for text-conditioned diffusion models that focuses on guiding text prompt embeddings rather than controlling the generative process. Our method constructs a semantic boundary that maximally distinguishes between safe and unsafe content. We then project potentially unsafe embeddings toward the safe region while preserving the original semantics and maintaining the diffusion model’s generative capabilities. This approach offers three key advantages: efficiency, effectiveness, and versatility. By operating at the prompt embedding level, our method eliminates the need for computationally intensive model re-training while preserving the original semantics. Moreover, by preventing unsafe content at the earliest stage of generation, we can block the formation of unsafe latent representations more directly and reliably. Lastly, our lightweight approach can be easily trained and applied to various concept removal tasks.
@@
We benchmark our prototype SteerDiff against state-of-the-art concept removal techniques, including Erased Stable Diffusion (ESD) (Gandikota et al., 2023) and Safe Latent Diffusion (SLD) (Schramowski et al., 2023; Rando et al., 2022), for removing inappropriate content. Experimental results demonstrate that SteerDiff significantly reduces inappropriate content generation while preserving image quality and semantic fidelity. Furthermore, we evaluate its robustness by defending against red-teaming methods such as P4D and SneakyPrompt, demonstrating the effectiveness of our approach in mitigating various forms of adversarial attacks on text-to-image generation systems.
2 Related work
2.1 Text-to-Image Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022) are a type of probabilistic generative model that learns a data distribution by gradually transforming a simple distribution into a complex target distribution. Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) model the data generation process as a sequence of denoising steps. Given an image, DDPMs progressively add noise sampled from Gaussian distribution to generate an intermediate noisy image $\mathbf{x}_{t}$ at each time step $t$ called forward diffusion steps. The noisy image $\mathbf{x}_{t}$ can be expressed in closed form as a function of the original image $\mathbf{x}_{0}$, the time step $t$, and noise $\mathbf{\epsilon}$ sampled from a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$. The model is then trained using the reverse process, where the objective is to learn a model parameterized by $\theta$ that predicts $\mathbf{\epsilon}$. This objective is defined as:
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\epsilon\theta(\mathbf{x}_{t},t)\|^{2}_{2}\right].$
2.2 Towards Safe Image Generation
Current approaches to prevent undesirable images from generation generally follow two main strategies. The first involves removing undesired images from the training set, such as excluding all human figures (Nichol et al., 2021) or selectively omitting specific undesirable image categories in the dataset (Schuhmann et al., 2022; DAL, 2022; Rombach & Esser, 2022). However, this approach is costly as it necessitates retraining the model, and removing certain data classes often degrades the overall quality of the generated images (O’Connor, 2022). The second strategy is post-hoc, includes using blacklists for blocking unsafe concepts (DAL, 2022; mid, 2024; Leo, 2023; Markov et al., 2023), editing images after generation, or fine-tuning diffusion models to guide the inference process (Schramowski et al., 2023). Although blacklists are straightforward to implement, they can be easily bypassed. Image editing and diffusion process manipulation methods are effective but still require image synthesis, adding computational overhead. Other approaches (Park et al., 2024; Zhang et al., 2024) use inpainting to mask unsafe content or attempt to unlearn inappropriate concepts either within the diffusion model (Zheng & Yeh, 2023) or the textual encoder (Poppi et al., 2023). These methods involve expensive fine-tuning, while SteerDiff offers an intermediary solution that operates between the input prompt and the diffusion model without additional training. Similarly, Latent Guard (Liu et al., 2024) identifies inappropriate concepts prior to diffusion by learning a latent space on top of the T2I model’s text encoder. However, while Latent Guard is capable of detecting inappropriate concepts, it lacks the ability to remove them. In this work, we compare the naive blacklist method commonly used in commercial platforms, Erased Stable Diffusion (Gandikota et al., 2023), the state-of-the-art model-editing approach, and Safe Latent Diffusion (Schramowski et al., 2023), a guidance-based model. Our focus is to introduce a new approach that identifies and manipulates inappropriate concepts in the text embedding space to ensure safer image generation.
3 Methodology
@@
Determining what constitutes inappropriate imagery is highly subjective and varies based on context, setting, cultural and social predispositions, and individual factors. Additionally, Qu et al. (2023) observes that unsafe content can be shared through memes. In the context of T2I models, these new concepts are embedded within the words of a prompt. Consequently, adding new concepts to block T2I generation without retraining the diffusion model is impractical. To overcome these limitations, we formalize the problem by identifying unsafe concepts in text embedding space and then projecting potentially unsafe embeddings toward safe regions. This projection preserves the original semantics while maintaining the diffusion model’s generative capabilities. Our approach allows us to define blacklisted concepts at test time, enabling greater flexibility.
Figure 1: Overview of data collection and training process for SteerDiff: We begin by defining a set of unsafe concepts, $C_{\text{unsafe}}$ (yellow block). Next, we use a LLM to generate related unsafe (red block on the left) and safe (green block on the left) phrases based on each concept $c\in C_{\text{unsafe}}$. These phrases are then encoded using a pre-trained text encoder to extract embedding features, which are used to train both the identifier and steering model.
@@
In subsection 3.1, we find a set of inappropriate concepts based on established work (Schramowski et al., 2023) and describe the process of collecting and generating training data using a large language model as illustrated in Figure 1. As shown in the overview in 2(a), we next explain how inappropriate concepts are identified within text embeddings (subsection 3.2), followed by our steering approach to mitigate the generation of inappropriate content (subsection 3.3). Finally, we outline how the framework operates during inference to identify and steer text prompts associated with unsafe concepts (subsection 3.4).
3.1 Training Data Collection
@@
[…]
@@
It is important to note that the definition of inappropriateness is not restricted to these seven categories, as the boundaries of appropriateness vary across cultures and evolve over time. In this study, however, we limit our scope to images that display clear and tangible acts of inappropriate behavior. Although we limit our scope to the current $C_{\text{unsafe}}$, our framework can be extended beyond these categories, as discussed in section 5.
3.1.1 Identifier Dataset
@@
As mentioned previously, we aim to detect inappropriate concepts in prompts to avoid inappropriate content generation. The first step in our pipeline is to construct the dataset with unsafe terms capturing the concepts from $C_{\text{unsafe}}$. To achieve this, we start by collecting open-sourced blacklisted phrases. Midjourney (mid, 2024) employs a blacklist of words and phrases that includes phrases associated with violence, hate speech, explicit sexual content, illegal activities, and other categories considered inappropriate by the platform. Additionally, Latent Guard (Liu et al., 2024) introduces the CoPro dataset, which includes safe and unsafe prompts centered around blacklisted concepts. We build our blacklist based on Midjourney and CoPro dataset.
@@
Although we can collect numerous NSFW terms from open-source datasets, these datasets may be imbalanced and could lack some categories defined in $C_{\text{unsafe}}$. For example, the number of shocking images or illegal activity is lower compared to images in other categories. To tackle this, we leverage an LLM to generate related terms $t_{c}$ centered around one sampled concept $c$ in the blacklist of $C_{\text{unsafe}}$ as illustrated in Figure 1, similarly to Hammoud et al. (2024) and Liu et al. (2024). This allows us to create a set $T_{\text{unsafe}}=\{t_{c}|c\in C_{\text{unsafe}}\}$. The unsafe terms in $T_{\text{unsafe}}$ simulate typical unsafe terms that a malicious user may input. In addition to the unsafe terms, we combine randomly selected 500 prompts (Blue block on the left) from the COCO 30k dataset (Lin et al., 2014) and the following mentioned paired safe phrases dataset (Green block on the left) to serve as our safe prompt dataset. All phrases within these prompts are considered safe.
3.1.2 Steer Model Dataset
@@
In our subsequent steering transformation training procedure, we synthesize additional safe terms to steer unsafe embeddings toward safe ones. The core idea is to associate each unsafe term $t_{c}\in T_{\text{unsafe}}$ with a corresponding safe term ${t_{c}}^{\prime}$ of similar meanings, allowing us to convert unsafe concepts into safe alternatives while preserving the original semantic intent of the prompt. For example, consider the term “killed” in the prompt “A man got killed.” which represents a violent visual scene linked to the concept of “violence”. We use an LLM to eliminate unsafe concepts in the input term $t_{c}$. In this case, a possible safe term ${t_{c}}^{\prime}$ would be “saved”, transforming the meaning of the prompt to something like “A man got saved”. By processing all elements in $T_{\text{unsafe}}$, we generate paired safe phrases dataset $T_{\text{safe}}$, comprising $M$ safe terms ${t_{c}}^{\prime}\in T_{\text{safe}}$. For more details of the synthesized dataset, please refer to subsection A.2.
3.2 Inappropriate Concepts Identifier
@@
To ensure the generation of safe and appropriate content, we detect and mitigate undesirable concepts within user prompts before they are processed by the text-to-image diffusion model. The core idea is that the prompt embeddings can be leveraged to represent individual terms or phrases, enabling precise identification of inappropriate content.
This task can be framed as classifying phrases as either appropriate or inappropriate. We employ a lightweight multi-layer perceptron model (MLP) to classify the embeddings of individual terms. Intuitively, we want to identify the phrase that precisely includes the inappropriate concept. For example, in the prompt “a man got shot.” only the term “shot” relates to the concept of violence. Additionally, the phrase “screw you” is related to the inappropriate concept “hate”, but neither “screw” nor “you” is related to the concept “hate”. To address such cases, we utilize a sliding window technique to identify single terms and phrases that may collectively express inappropriate concepts.
@@
As shown in Figure 1, the safe and unsafe phrases are first embedded by a text encoder and then used to train the identifier. We expect the identifier to detect the unsafe concepts defined in Equation 1. Let $T_{\text{unsafe}}$ represent the set of unsafe phrases. Specifically, the safe phrases set is composed of two subsets: $T_{\text{safe}}$ (synthetic set) and $T_{\text{safe}’}$ (collected from open-source resources). The MLP outputs the probability of a given phrase belonging to unsafe or safe. Denote the input phrase embeddings as $E_{i}$, the MLP output as $\hat{y}_{i}$, and the true label as $y_{i}\in\{0,1\}$, where $y_{i}=1$ indicates an unsafe phrase and $y_{i}=0$ indicates a safe phrase. The classification loss $\mathcal{L}$ is defined as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right)$ (2)
@@
Since embeddings with similar semantics have closer distances in the embedding space (Mikolov et al., 2013; Radford et al., 2021), we expect the unsafe embeddings to be aggregated. As demonstrated in 2(b), we observe that SteerDiff successfully learns to distinguish between safe and unsafe phrases, with the two categories being well-separated after applying t-SNE dimensional reduction.
(a) A T2I generator without a safety mechanism (top) can generate inappropriate content. We propose SteerDiff (bottom), a safety method designed to identify and steer inappropriate concepts toward producing safe images.
Figure 2: Overview of SteerDiff (left). SteerDiff learns to distinguish safe and unsafe phrases (right).
3.3 Steering Toward Safe Content

### v2

@@
Abstract.
Text-to-image (T2I) diffusion models have drawn attention for their ability to generate high-quality images with precise text alignment. However, these models can also be misused to produce harmful or inappropriate content. Existing safety measures, which typically rely on text classifiers or ControlNet-like approaches, are often insufficient. Text classifiers rely on large-scale labeled datasets and can be easily bypassed by rephrasing. As diffusion models continue to scale, fine-tuning safeguards becomes increasingly challenging and lacks flexibility. Recent red-teaming attack researches further underscore the need for a new paradigm to prevent the generation of inappropriate content. In this paper, we introduce SteerDiff, a lightweight adapter module designed to act as an intermediary between user input and the diffusion model, ensuring that generated images adhere to ethical and safety standards with little to no impact on usability. SteerDiff identifies and manipulates inappropriate concepts within the text embedding space to guide the model away from harmful outputs. We conducted extensive experiments across multiple concept unlearning tasks to evaluate the effectiveness of our approach. Furthermore, we benchmark SteerDiff against multiple red-teaming strategies to assess its robustness. Finally, we explore the potential of SteerDiff for artist-style removal tasks, demonstrating its versatility in text-conditioned image generation.
@@ Abstract.
Diffusion Model, Text-to-Image, Concept Unlearning
copyright: noneconference: ; ; $c$s: Computing methodologies Computer visionccs: Security and privacy Social aspects of security and privacy
1. Introduction
@@
Text-to-image (T2I) diffusion models have attracted attention for their out-of-the-box functionality and the superior quality of their generated images. Using models like Stable Diffusion (Rombach et al., 2022, 2021) or DALL-E (Ramesh et al., 2022), users can use simple natural language descriptions as input to generate high-quality images with precise text alignment. This capability is largely contributed by pre-trained language models (Hammoud et al., 2024; Jin et al., 2020; Dosovitskiy, 2020) that learn and reflect the underlying syntax and semantics, as well as by extensive multimodal training datasets that encompass a wide range of text-to-image aligned content. However, these training methods also introduce the risk of generating inappropriate content with unsafe concepts such as sex, violence, or illegal activity. Thus, preventing the generation of inappropriate images is both critical and urgent.
To tackle these challenges, recent research has integrated several safety strategies to prevent inappropriate content generation. Model editing methods (Gandikota et al., 2023; Orgad et al., 2023; Gandikota et al., 2024) identify and prune weights within the diffusion model to remove harmful concepts, but they often degrade output quality, thus harming the utility of the models (Rombach et al., 2022). Unlearning methods(Zhang et al., 2024b; Huang et al., 2024) finetune models to erase harmful content, but as model sizes grow, this becomes increasingly computationally expensive and less practical(Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022). A practical alternative lies in training-free, filtering-based methods, which aim to exclude unsafe concepts directly from input prompts without modifying the underlying model. However, existing approaches in this category (Liu et al., 2024; Schuhmann et al., 2022; Schramowski et al., 2023; Rando et al., 2022) can be easily passed by text-based attacks (Li et al., 2018; Jin et al., 2020; Garg and Ramakrishnan, 2020; Maus et al., 2023) through mislead the classification mechanisms in the filter by rephrasing, e.g.“A photo of a billboard showing a naked man.” into “A photo of a billboard showing an LGBT man in an explicit position”.
@@
Although the aforementioned safety mechanisms have shown effectiveness according to their respective evaluation schemes, recent red-teaming studies demonstrate their potential flaws (Zhang et al., 2023; Qu et al., 2023). Chin et al. (2023) show that approximately half of the prompts, which were originally blocked by existing safety mechanisms, can be manipulated by their Prompting4Debugging (P4D) to become problematic. Similarly, Unsafe Diffusion (Qu et al., 2023) finds that 14.56% of generated images across four state-of-the-art T2I models and their prompt datasets were unsafe, underscoring the vulnerability of these models to generating harmful content. Moreover, black-box jailbreaking approaches Jailbreak Prompt Attack (JPA) and SneakyPrompt (Ma et al., 2024; Yang et al., 2024b) successfully attack both online services and offline T2I models with the current safety mechanisms.
@@
In this work, we propose SteerDiff, a novel lightweight adaptor model for text-conditioned diffusion models to ensure safe generation without altering model weights. As illustrated in Figure 1, SteerDiff first constructs a semantic boundary distinguishing between safe and unsafe content. Then, it projects potentially unsafe embeddings toward the safe region while preserving the original semantics and maintaining the diffusion model’s generative capabilities. This approach offers three key advantages: efficiency, effectiveness, and versatility. By operating at the prompt embedding level, our method eliminates the need for computationally intensive model re-training while preserving the original semantics. Moreover, by filtering unsafe content early in the generation process, SteerDiff prevents harmful latent representations and achieves state-of-the-art robustness in our red-teaming experiments.
Figure 1. Overview of SteerDiff, a safety method designed to identify and steer inappropriate concepts toward producing safe images. The input prompt is first embedded by a text encoder. The identifier then checks if the prompt contains any inappropriate concepts $c^{\prime}\in C_{\text{unsafe}}$. If detected, a linear transformation is applied to the prompt’s embedding to steer it toward safer content. This transformation adjusts the embedding space while preserving the semantics of the original prompt. Once transformed, the modified embedding is passed through the diffusion model to generate safe images.
@@
We benchmark SteerDiff against state-of-the-art concept removal techniques, including Erased Stable Diffusion (ESD) (Gandikota et al., 2023), SPM (Lyu et al., 2024) and Safe Latent Diffusion (SLD) (Schramowski et al., 2023; Rando et al., 2022), for removing inappropriate content. Experimental results demonstrate that SteerDiff significantly reduces inappropriate content generation while preserving image quality and semantic fidelity. Furthermore, we evaluate its robustness by defending against red-teaming methods, including Prompt4Debugging, Ring-A-Bell, MMA-Diffusion, and SneakyPrompt(Chin et al., 2023; Tsai et al., 2023; Yang et al., 2024a, b), demonstrating the effectiveness of our approach in mitigating various forms of adversarial attacks on text-to-image generation systems. Lastly, we evaluate the potential of SteerDiff for artist-style removal tasks, highlighting its versatility in text-conditioned image generation. In summary, our main contributions are as follows:
  • We propose a novel lightweight adaptor model SteerDiff for T2I diffusion models to ensure safe image generation without altering the model weights.
  • We comprehensively evaluate SteerDiff across different scenarios and red-teaming attacks, achieving state-of-the-art robustness, semantic consistency, and image quality.
  • SteerDiff shows versatility for other tasks, such as artist-style removal, underscoring its broad applicability in text-conditioned image generation scenarios.
2. Methodology
@@
Determining what constitutes inappropriate imagery is highly subjective and varies based on context, setting, cultural and social predispositions, and individual factors. Additionally, (Qu et al., 2023) observes that unsafe content can be shared through memes. In the context of T2I models, these new concepts are embedded within the words of a prompt. Consequently, adding new concepts to block T2I generation without retraining the diffusion model is impractical. To overcome these limitations, we formalize the problem by identifying unsafe concepts in text embedding space and then projecting potentially unsafe embeddings toward safe regions. This projection preserves the original semantics while maintaining the diffusion model’s generative capabilities. Our approach allows us to define blacklisted concepts at test time, enabling greater flexibility.
@@
In subsection 2.1, we find a set of inappropriate concepts based on established work (Schramowski et al., 2023) and describe the process of collecting and generating training data using a large language model as illustrated in 2(a). As shown in the overview in Figure 1, we next explain how inappropriate concepts are identified within text embeddings (subsection 2.2), followed by our steering approach to mitigate the generation of inappropriate content (subsection 2.3). Finally, we outline how the framework operates during inference to identify and steer text prompts associated with unsafe concepts (subsection 2.4).
2.1. Training Data Collection
@@
(1) $C_{\text{unsafe}}=$ $\{\text{hate, harassment, violence, self-harm, sexual }$
$\text{content, shocking images, illegal activity}\}$
@@
It is important to note that the definition of inappropriateness is not restricted to these seven categories, as the boundaries of appropriateness vary across cultures and evolve over time. In this study, we limit our scope to images displaying clear and tangible inappropriate behavior acts. Although we limit our scope to the current $C_{\text{unsafe}}$, our framework can be extended beyond this task, as discussed in subsection 3.6.
(a) Overview of data collection and training process: We begin with a defined set of unsafe concepts, $C_{\text{unsafe}}$ (yellow block). Next, we use an LLM to generate related unsafe (red block on the left) and safe (green block on the left) phrases based on each concept $c\in C_{\text{unsafe}}$. Additionally, we incorporated a set of safe phrases from open sources (blue block) to serve as a part of safe embeddings. These phrases are then encoded to train the identifier and steering model.
Figure 2. Overview of SteerDiff (left). SteerDiff learns to distinguish safe and unsafe phrases (right).
Identifier Dataset
@@
As mentioned previously, we aim to detect inappropriate concepts in prompts to avoid inappropriate content generation. The first step in our pipeline is to construct the dataset with unsafe terms capturing the concepts of $C_{\text{unsafe}}$. To achieve this, we start by collecting open-sourced blacklisted phrases. Midjourney (mid, 2024) employs a blacklist of words and phrases that includes phrases associated with violence, hate speech, explicit sexual content, illegal activities, and other categories considered inappropriate by the platform. Additionally, Latent Guard (Liu et al., 2024) introduces the CoPro dataset, which includes safe and unsafe prompts centered around blacklisted concepts. We build our blacklist based on Midjourney and CoPro dataset.
@@
Although we can collect numerous NSFW terms from open-source datasets, these datasets may be imbalanced and could lack some categories defined in $C_{\text{unsafe}}$. For instance, the amount of data in categories involving shocking or illegal activities is lower than in other categories. To tackle this, we leverage an LLM to generate related terms $t_{c}$ centered around one sampled concept $c$ in the blacklist of $C_{\text{unsafe}}$ as illustrated in 2(a), similarly to Hammoud et al. (2024) and Liu et al. (2024). This allows us to create a set $T_{\text{unsafe}}=\{\;t_{c}\;|\;c\in C_{\text{unsafe}}\}$. The unsafe terms in $T_{\text{unsafe}}$ (Red block on the left) simulate typical unsafe terms that a malicious user may input. In addition to the unsafe terms, we combine randomly selected 500 prompts (Blue block on the left) from the COCO 30k dataset (Lin et al., 2014) and the following mentioned paired safe phrases dataset (Green block on the left) to serve as our safe prompt dataset. All phrases within these prompts are considered safe.
Steer Model Dataset
@@
In our subsequent steering transformation training procedure, we synthesize additional safe terms to steer unsafe embeddings toward safe ones. The core idea is to associate each unsafe term $t_{c}\in T_{\text{unsafe}}$ with a corresponding safe term $t_{c}$$^{\prime}$ of similar meanings, allowing us to convert unsafe concepts into safe alternatives while preserving the original semantic intent of the prompt. For example, consider the term “killed” in the prompt “A man got killed.” which represents a violent visual scene linked to the concept of “violence”. As demonstrated in 2(a), we use GPT-4 (OpenAI, 2023) to eliminate unsafe concepts in the input term $t_{c}$. In this case, a possible safe term $t_{c}$$^{\prime}$ would be “saved”, transforming the meaning of the prompt to something like “A man got saved”. By processing all elements in $T_{\text{unsafe}}$, we generate paired safe phrases dataset $T_{\text{safe}}$ (Green block on the left), comprising $M$ safe terms ${t_{c}}^{\prime}\in T_{\text{safe}}$. Statistics of the dataset are listed in supplementary materials.
2.2. Inappropriate Concepts Identifier
@@
To ensure the generated images adhere to safety guidelines, we detect and mitigate undesirable concepts within user prompts before they are processed by the text-to-image diffusion model. The core idea is that the prompt embeddings can be leveraged to represent individual terms or phrases, enabling precise identification of inappropriate content.
@@
As shown in 2(a), both safe and unsafe phrases are first embedded using a text encoder and then utilized to train the identifier. The goal is for the identifier to accurately detect unsafe concepts as defined in Equation 1. Let $T_{\text{unsafe}}$ denote the set of unsafe phrases. Specifically, the set of safe phrases consists of two subsets: $T_{\text{safe}}$, containing synthetic phrases, and $T_{\text{safe}^{\prime}}$, comprising phrases collected from open-source resources.
The identifier’s task is to classify a given phrase as either unsafe or safe. Let the input phrase embeddings be denoted as $E_{i}$, the identifier’s output as $\hat{y}_{i}$, and the true label as $y_{i}\in\{0,1\}$, where $y_{i}=1$ indicates an unsafe phrase and $y_{i}=0$ denotes a safe phrase. The identifier is optimized using binary cross-entropy loss:
(2) $\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log\hat{y}_{i}+(1-y_{i})\cdot\log(1-\hat{y}_{i})\right]$
@@
Since embeddings with similar semantics are known to cluster in the embedding space (Mikolov et al., 2013; Radford et al., 2021), we expect unsafe embeddings to naturally aggregate; thus, SteerDiff can demonstrate a generalization capability. As demonstrated in 2(b), a t-SNE visualization shows a clear separation between safe and unsafe embedding spaces derived from MMA (Yang et al., 2024a) and COCO (Lin et al., 2014) datasets. Moreover, the experimental results in subsection 3.4 further confirm that SteerDiff maintains robust performance against various red-teaming attacks.
In practice, we employ a lightweight multi-layer perceptron model to classify the embeddings of individual terms. Intuitively, we want to identify the phrase that precisely includes the inappropriate concept. For example, in the prompt “a man got shot.” only the term “shot” relates to the concept of violence. Additionally, the phrase “screw you” is related to the inappropriate concept “hate”, but neither “screw” nor “you” is related to the concept “hate”. To address such cases, we utilize a sliding identification technique to identify single terms and phrases that may collectively express inappropriate concepts.
Algorithm 1 Sliding Identification
1:Initialize: window size $w$, position map $\mathcal{P}$, token length $L$, prompt vector set $\mathcal{V}$
2:for $i=1$ to $L$ do
3:  for $j=1$ to $w$ do
4:   $t\leftarrow\min(L,i+j)$ $\triangleright$ Set tail index
5:   Extract token embeddings $e[i:t]$
6:   Pad to fixed size and store in $\mathcal{V}$
7:   Map token span in $\mathcal{P}$
8:  end for
9:end for
10:Predict unsafe probabilities for $\mathcal{V}$ using identifier model
Sliding Identification
To detect inappropriate concepts with fine-grained precision, we introduce a sliding identification mechanism that captures distributed unsafe signals across contiguous token spans. This method systematically scans a prompt using a sliding window over token embeddings, allowing the model to identify not only individual offensive terms but also compound phrases whose semantic meaning emerges only when considered jointly. As illustrated in 1, for each starting token in the prompt, we construct multiple token spans of varying lengths up to a predefined window size $w$. Each span is embedded, padded to a fixed size, and evaluated by the identifier to assess if it contains unsafe content.
2.3. Steering Toward Safe Content